OpenStudy (anonymous):

Find any critical numbers of the function. (sin(x))^2 + cos(x) ) 0 < x < 2π

zepdrix (zepdrix):

Were you able to find the derivative function?

OpenStudy (anonymous):

not sure if its correct : 2 (sin x) (cos x) - sin x ?

zepdrix (zepdrix):

Yah looks good :)

OpenStudy (anonymous):

ok then what? hehe

zepdrix (zepdrix):

Critical points exist where lines tangent to our function are horizontal. Or in other words, when our derivative function is zero. \[\large 2\sin x \cos x - \sin x=0\] So we let our function equal zero like this, and solve for x.

zepdrix (zepdrix):

If we factor a sin x out of each term, it gives us,\[\large \sin x(2\cos x-1)=0\]right? :o

OpenStudy (anonymous):

yupp i got it now i know what to do ! thanks !

zepdrix (zepdrix):

ok cool :)

OpenStudy (anonymous):

thanks!